# Residual Kolmogorov-Arnold Network for Enhanced Deep Learning

### v3

@@ 4.6 Object Detection and Instance Segmentation
We apply Mask R-CNN for object detection and instance segmentation [16] on MS COCO 2017. We then use our proposed RKANet-50 to compare with other enhancement methods (e.g. ECA-Net). Models are trained for 12 epochs using pre-trained weights from ImageNet-1k. We apply SGD with a weight decay of $\text{10}^{-4}$, batch size of 8, and base learning rate of 0.01 that decays by a factor of 10 at epoch 9 and 12. For data augmentation, we apply basic horizontal flipping ($p$ = 0.5), random scaling ($\leq$ 10%), color jittering.
@@ Appendix A Implementation Details
ConvNeXt-S [37] AdamW 0.05 0.004
MaxViT-T [52] AdamW 0.05 0.003
EfficientViT-L2 [51] AdamW 0.05 0.001
FastViT-SA24 [53] AdamW 0.05 0.001
CoAtNet-1 [7] AdamW 0.1 0.001
MobileNetV4-L [43] SGD 1e-4 0.05
EfficientNetV2-M [51] SGD 1e-4 0.05
@@ Appendix B Inspiration
Wide ResNet [62] has demonstrated the potential of wider networks. For example, WRN-50-2 has a reduction ratio of 2 in the bottleneck layers (128, 256, 512, 1024), whereas ResNet-101 retains a reduction ratio of 4. In comparison, WRN-50-2 outperforms the latter with half the network depth in most tasks, such as CIFAR-100 and ImageNet-1k, showing that ultra-deep network is not the only solution.
@@
Figure 4: FLOPs and throughput compared between RBF-based and Chebyshev polynomial-based RKAN-augmented models on the ILSVRC-2012 ImageNet dataset [8] of resolution $\times224$.
@@ Appendix C Computational Metrics
The total number of model parameters for a KAN-based convolutional layer is primarily determined by the degree ($d$) of Chebyshev polynomials in comparison to a standard convolutional layer, expressed as:
$KAN_{\text{p}}=C_{in}\times C_{out}\times k_{h}\times k_{w}\times(d+1)$ (9)
$C_{in}$ and $C_{out}$ denote the number of input and output channels, respectively, while $k_{h}\times k_{w}$ is the kernel size (height × width). As shown in Table˜9, model parameters for all RKAN-augmented ResNet architectures that have the bottleneck structure ($C_{in}=1024$, $C_{out}=2048$) with $r=2$ are increased by a fixed amount of 1.58 million. For example, RKANet-152 retains 2.7% more parameters and 1.4% more FLOPs than baseline ResNet-152. In contrast, RKANet-152 reduces the throughput by 12.9% compared to the baseline, which is hardly reflected in the calculation of either metrics above.
Figure 5: KAN-based convolutional layer implemented using Chebyshev polynomials.
@@
Appendix D More Details on KAN Convolution
We demonstrate the process of how features are processed within a single KAN convolutional layer, shown in Figure˜5. The input tensor should match the output spatial dimension ($H_{in}\times W_{in}$) of the first $\times1$ bottleneck layer while the output tensor needs to match the input spatial dimension ($H_{out}\times W_{out}$) of the second bottleneck layer and the main network path output. The channel-wise dimension usually remains unchanged and is regulated by the bottleneck layers for efficiency instead. Algorithm˜1 provides a more detailed description (pseudocode) of the convolution operation.
The polynomial-based KAN layer, represented by the “Chebyshev Expansion" in Figure˜5 is primarily inspired by the original spline-based KAN implementation where the activation function $\phi(x)$ is a linear combination of the SiLU activation $\mathrm{silu}(x)$ and the spline function [39], in which $N$ denotes the number of splines, $B_{i}(x)$ are the B-spline basis functions, and $c_{i}$ are the trainable coefficients:

### v4

@@ 4.6 Object Detection and Instance Segmentation
We apply Mask R-CNN for object detection and instance segmentation [he2017mask] on MS COCO 2017. We then use our proposed RKANet-50 to compare with other enhancement methods (e.g. ECA-Net). Models are trained for 12 epochs using pre-trained weights from ImageNet-1k. We apply SGD with a weight decay of $\text{10}^{-4}$, batch size of 8, and base learning rate of 0.01 that decays by a factor of 10 at epoch 9 and 12. For data augmentation, we apply basic horizontal flipping ($p$ = 0.5), random scaling ($\leq$ 10%), color jittering.
@@ Appendix A Implementation Details
ConvNeXt-S [liu2022convnet] AdamW 0.05 0.004
MaxViT-T [tu2022maxvit] AdamW 0.05 0.003
EfficientViT-L2 [tan2021efficientnetv2] AdamW 0.05 0.001
FastViT-MA36 [vasu2023fastvit] AdamW 0.05 0.001
CoAtNet-1 [dai2021coatnet] AdamW 0.1 0.001
MobileNetV4-L [qin2024mobilenetv4] SGD 1e-4 0.05
EfficientNetV2-M [tan2021efficientnetv2] SGD 1e-4 0.05
@@ Appendix B Inspiration
Wide ResNet [zagoruyko2016wide] has demonstrated the potential of wider networks. For example, WRN-50-2 has a reduction ratio of 2 in the bottleneck layers (128, 256, 512, 1024), whereas ResNet-101 retains a reduction ratio of 4. In comparison, WRN-50-2 outperforms the latter with half the network depth in most tasks, such as CIFAR-100 and ImageNet-1k, showing that ultra-deep network is not the only solution.
@@
Figure 4: FLOPs and throughput compared between RBF-based and Chebyshev polynomial-based RKAN-augmented models on the ILSVRC-2012 ImageNet dataset [deng2009imagenet] of resolution $\times224$.
@@ Appendix C Computational Metrics
The total number of model parameters for our KAN-based convolutional layer is primarily determined by the degree ($d$) of Chebyshev polynomials in comparison to a standard depthwise separable convolutional layer, expressed as:
$KAN_{\text{p, DWC}}=C_{in}\times k_{h}\times k_{w}\times(d+1)$ (9)
$C_{in}$ denotes the input channels, while $k_{h}\times k_{w}$ is the kernel size (height × width). As shown in Table˜9, model parameters for all RKAN-augmented ResNet architectures that have the bottleneck structure ($C_{in}=1024$, $C_{out}=2048$) are increased by a fixed amount of 1.58 million. For example, RKANet-152 has 2.7% more model parameters and 1.4% more FLOPs than baseline ResNet-152. However, RKANet-152 reduces the throughput by 12.9% compared to the baseline, which is hardly reflected in the calculation of either metrics above.
Figure 5: KAN-based convolutional layer implemented using Chebyshev polynomials, applied depthwise (independently to each channel).
@@
Appendix D Details on Depthwise KAN Convolution
We demonstrate the process of how features are processed within a single KAN convolutional layer, shown in Figure˜5. The input tensor should match the output spatial dimension ($H_{in}\times W_{in}$) of the first $\times1$ bottleneck layer while the output tensor needs to match the input spatial dimension ($H_{out}\times W_{out}$) of the second bottleneck layer and the main network path output. The channel-wise dimension remains unchanged (due to depthwise convolution) and is regulated by the bottleneck layers for efficiency instead. Algorithm˜1 provides a more detailed description of the operation.
The polynomial-based KAN layer, represented by the “Chebyshev Expansion" in Figure˜5 is primarily inspired by the original spline-based KAN implementation where the activation function $\phi(x)$ is a linear combination of the SiLU activation $\mathrm{silu}(x)$ and the spline function [liu2023kolmogorov], in which $N$ denotes the number of splines, $B_{i}(x)$ are the B-spline basis functions, and $c_{i}$ are the trainable coefficients: